OpenStudy (nene101):

F(x)=-11x^2 domain and range ?

OpenStudy (thatonegirl_):

if you plug in any number for x, will you get always get an answer back?

OpenStudy (nene101):

Yes

OpenStudy (nene101):

So -infinity ?

OpenStudy (thatonegirl_):

so that means your domain (number you plug in) can be any real number. so you can write that as all real numbers or as \[(-\infty,\infty)\] depending on what you have been doing in class

OpenStudy (thatonegirl_):

And what kind of number will you get back for any x value? consider plugging in both negative x's and positive ones. What will your y value be?

OpenStudy (nene101):

So (- infinity, + infinity)

OpenStudy (nene101):

For both domain and range ?

OpenStudy (thatonegirl_):

For the range? the values you get back?

OpenStudy (thatonegirl_):

Well domain you had correct. Any and all numbers from negative infinity to positive infinity. But what happens when you plug in a negative number for x? What is your y? positive or negative?

OpenStudy (nene101):

Postive

OpenStudy (nene101):

Positive

OpenStudy (nene101):

(0,infinity)

OpenStudy (thatonegirl_):

Hmm.. \[f(x)=-11x^2\] \[f(x)=-11(-10)^2\] \[f(x)=-11(100)\] \[f(x)=-1100\]

OpenStudy (nene101):

Ohh I see, negative

OpenStudy (thatonegirl_):

So a negative answer right? You can plug in any negative number and you will get a negative answer. How about if you plug in a positive x?? Will your answer be positive or negative?

OpenStudy (nene101):

(-infinity , 0]

OpenStudy (thatonegirl_):

Yes perfect!

OpenStudy (thatonegirl_):

So now you have your domain: \[(-\infty,\infty)\] and the range: \[(-\infty,0 ]\]

OpenStudy (thatonegirl_):

Make sense?

OpenStudy (nene101):

Yess thanks so much